OpenStudy (anonymous):

Find the lateral area of a right circular cone whose radius is 6 feet and height is 8 feet.

OpenStudy (anonymous):

lateral are - pi*r*s where s = slant height s = sqrt(8^2 + 6^2) = 10 snalt area = pi*6*10 sq ft = 188.5 sq ft